### v1

# On completeness of Foliated structures, and Null Killing fields

Abstract.
We consider a compact manifold $(M,\mathfrak{F})$ with a foliation $\mathfrak{F}$, and a smooth affine connection $\nabla$ on the tangent bundle of the foliation $T\mathfrak{F}$. We introduce and study a foliated completeness problem. Namely, under which conditions on $\nabla$ the leaves are complete? We consider different natural geometric settings: the first one is the case of a totally geodesic lightlike foliation of a compact Lorentzian manifold, and the second one is the case where the leaves have particular affine structures. In the first case, we characterize the completeness, and obtain in particular that if a compact Lorentzian manifold admits a null Killing field $V$ such that the distribution orthogonal to $V$ is integrable, then it defines a (totally geodesic) foliation with complete leaves. In the second case, we give a completeness result for a specific affine structure called “the unimodular affine lightlike geometry”, and characterize the completeness for a natural relaxation of the geometry. On the other hand, we study the global completeness of a compact Lorentzian manifold in the presence of a null Killing field. We give two non-complete examples, starting from dimension $3$: one is a locally homogeneous manifold, and the other is a $3$D example where the Killing field dynamics is equicontinuous.
1. Introduction
Let $(M,\nabla)$ be a (connected) compact smooth manifold equipped with a smooth affine connection on the tangent bundle $TM$. A natural question to ask is: under which assumption on the couple $(M,\nabla)$ the manifold is geodesically complete (i.e. every geodesic is defined on $\mathbb{R}$)? Historically this started for $M$ having a smooth Riemannian metric $g$ and $\nabla$ being the associated Levi-Civita connection. Such a couple is always geodesically complete. However, for a non-Riemannian metric, the analogous statement is not true in general. Thus, the study is more rich in this direction, especially in Lorentzian signature, by providing sufficient conditions (curvature conditions, existence of Killing (or parallel) fields,..) for a compact Lorentzian manifold to be geodesically complete [6, 21, 16, 17, 18].
1.1. Leafwise completeness.
In this paper we introduce a foliated version of this completeness problem. Namely, we consider a triple $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$, where $M$ is a compact manifold, $\mathfrak{F}$ a foliation of $M$, and $\nabla_{\mathfrak{F}}$ a smooth affine connection defined on $T\mathfrak{F}$. The completeness problem above can be seen as the codimension 00 case in this setting. We formulate the problem as follows
Problem 1.1.
Let $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$ be a triple as above. Are there natural (geometric) settings in which the leaves of $\mathfrak{F}$ are geodesically complete?
Of course, when the connection on a leaf of the foliation is the Levi-Civita connection of some Riemannian metric on the leaf, the geodesics contained in this leaf are complete. We refer to this as ‘a Riemannian situation’: this is an elementary situation from the point of view of the completeness problem. So we are interested in non-Riemannian situations. We consider in the following question a natural and interesting non-Riemannian situation:
Question 1.2 (Foliated lightlike geometry: leafwise completeness).
Let $(M,\mathfrak{F},\nabla_{\mathfrak{F}})$ be a triple as above. Assume that we have a degenerate Riemannian metric defined on the tangent bundle $T\mathfrak{F}$, such that on each leaf, the induced degenerate Riemannian metric is parallel with respect to the existing connection. Assume further that the kernel of the metric has dimension $1$, and denote by $\mathcal{N}$ the $1$-dimensional foliation it defines. Observe that the leaves of $\mathcal{N}$ are totally geodesic (indeed, the degenerate Riemannian metric on each leaf is parallel, so its kernel is invariant by parallel transport). Is it true that the leaves are complete if and only if the geodesics tangent to $\mathcal{N}$ are complete?
We do not give an answer in the full generality of Question 1.2, but this question contains many natural situations, which we study in this paper. One of them is that of a totally geodesic lightlike (i.e. the metric induced on each leaf is Riemannian degenerate) foliation $\mathfrak{F}$ of a compact Lorentzian manifold. In this case, the connection on the tangent bundle $T\mathfrak{F}$ is the restriction to $T\mathfrak{F}$ of the Levi-Civita connection of the Lorentzian metric. Another natural and important situation is when $\nabla$ is a flat connection. In this case, the leaves have some specific affine structure. The latter case actually belongs to a more general situation arising from Problem 1.1: that of a foliation with a $(G,X)$-structure on the leaves. We introduce the notion of a $(G,X)$-foliation in Definition 2.3, and we focus on certain affine structures with interesting application in the study of completeness in Lorentzian (or semi-Riemannian) geometry. In this case, the geodesic completeness is equivalent to the completeness of the structure when the model is complete. We refer to Subsection 1.3 of the introduction for the treatment of the problem in this setting.
To summarize, we consider the two following situations
  • (A)
    A compact Lorentzian manifold with a totally geodesic lightlike foliation. This solves Question 1.2 when the connection is induced by an ambient Lorentzian Levi-Civita connection.
  • (B)
    A foliated compact manifold where the leaves have an affine structure in the sense of $(G,X)$-structures (see Definition 2.3). Observe that here, there is no assumption on existence of an ambient connection. The most important case of Item (B) we consider here is the case of the ‘affine lightlike geometry’ (see Subsection 1.3). The latter solves Question 1.2 when $\nabla$ is a flat connection.
For Item (A), we prove the following
Theorem 1.3.
Let $M$ be a compact Lorentzian manifold. Let $\mathfrak{F}$ be a codimension $1$ totally geodesic lightlike foliation of $M$. Then the leaves of $\mathfrak{F}$ are geodesically complete with respect to the induced connection if and only if the null geodesics in the leaves of $\mathfrak{F}$ are complete.
Theorem 1.3 contains the class of (compact) Brinkmann spacetimes. A Brinkmann spacetime is a Lorentzian manifold with a parallel null vector field $V$, so the codimension $1$ foliation $\mathfrak{F}$ tangent to $V^{\perp}$ is lightlike and totally geodesic. It is proved in [18] that a compact Brinkmann spacetime is complete. So this contains in particular a leafwise completeness result (of geodesics tangent to $\mathfrak{F}$). Theorem 1.3 can be seen as a generalization of the leafwise completeness to the situation where only a foliation is given, and this contains for instance the wider class of (compact) Kundt spacetimes (see [5, Proposition 2.1]).
As an application we get the following consequences
Corollary 1.4.
Let $(M,l)$ be a compact Lorentzian manifold with a parallel null line field $l$. The foliation tangent to $l^{\perp}$ is geodesically complete if and only if the (null) geodesics tangent to $l$ are complete.
In [17, Section 4.3, Example 3], the authors ask whether a compact Lorentzian manifold with a parallel null line field $l$ has a foliation (tangent to $l^{\perp}$) with complete leaves. The corollary above gives a necessary and sufficient condition for this to be true. Moreover, we give an example in which this condition does not hold, leading to incompleteness along the leaves (Example 3.5).
Corollary 1.5.
Let $(M,V)$ be a compact Lorentzian manifold with a null Killing field. Assume that the $V^{\perp}$-distribution is integrable (i.e. involutive). Then the foliation tangent to $V^{\perp}$ is geodesically complete.
1.2. On the geodesic completeness in presence of a null Killing field
For a compact Lorentzian manifold with a null Killing field, one could in fact consider the question of global completeness, i.e. of the whole manifold $M$. This question is not treated in the literature. When the metric admits a timelike Killing field $V$, it was observed in [21] that the intersection of the levels of the metric with the levels of $g(\cdot,V)$ are compact, which leads to completeness. However, this is no longer the case when $V$ is lightlike or spacelike. In fact, in the spacelike case, taking a product of a non-complete Lorentzian manifold with the circle (with a Riemannian metric) shows that existence of a spacelike Killing field does not imply completeness. But there is no known incomplete example in the lightlike case.
1.2.1. An incomplete example
We construct in this paper two examples of incomplete compact Lorentzian manifolds with a null Killing field. The first one is locally homogeneous, given by the quotient of a left invariant metric on $G:={\mathsf{Sol}}\times\mathbb{R}$. We state
Theorem 1.6.
There exist an incomplete left invariant Lorentzian metric on $G:={\mathsf{Sol}}\times\mathbb{R}$ with a central null Killing field $V$. Taking a quotient $M:=\Gamma\backslash G$, where $\Gamma$ is a cocompact lattice of $G$, defines an incomplete compact Lorentz manifold on which the flow of $V$ acts isometrically.
1.2.2. The case of dimension $3$
It is an easy fact that a compact Lorentzian surface with a null Killing field is complete (it is even flat). Our second incomplete example is more elaborate and is constructed in dimension $3$. Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$, with $V$ a lightlike Killing field. The first observation is that in this dimension, the geodesics orthogonal to $V$ are complete. For the global completeness question, it turns out that the situation is different, depending on the dynamics of $V$: when $V$ has complicated dynamics (non-equicontinuous), the manifold is complete; in the opposite situation (equicontinuous case), we construct an incomplete example (see paragraphs below).
I) Non-equicontinuous case
In this case, the closure of the flow of $V$ in the isometry group is non-compact. The global behaviour of $V$ is well understood [25]: Zeghib classifies all such flows on Lorentzian $3$-manifolds. In particular, he proves that $V^{\perp}$ is integrable and geodesic. And it follows from the classification that $M$ is complete [25, Remark 16.1]. Completeness for geodesics transversal to $V^{\perp}$ seems non-obvious in general, and the question is still open in higher dimension.
II) Equicontinuous case
The case where the flow of $V$ is equicontinuous is still completely unknown. In this case, the closure of the flow of $V$ in the isometry group is an abelian compact subgroup, hence a torus of some dimension $k$. We show that (Proposition 7.3) if $k\geq 3$, the manifold is just a flat $3$-torus. When $k=2$, we show completeness in some special situations, namely when the $\mathbb{T}^{2}$-action is free and the fibers of the $\mathbb{T}^{2}$-principal bundle have the same signature. Finally, we focus on the case where $k=1$. In this case, the action of $V$ is periodic, and when it is free, $M$ is an $\mathbb{S}^{1}$-bundle over a torus. The study of this situation raises interesting general questions on the relationship between the geodesic completeness and the dynamics of certain vector fields defined on the torus (we hope to come back to it in a future work).
Theorem 1.7.
There exist an incomplete compact $3$-dimensional Lorentzian manifold with a null Killing field.
The next subsection is devoted to the second part (B).
1.3. $(G,X)$-foliated completeness
Completeness of compact $(G,X)$-manifolds is a very rich and active topic of old and current research. It has been developed in many directions, especially in the case of affine structures, for instance [6, 9, 8, 3]. Completeness of compact $(G,X)$-manifolds is in general far from being easy. Here we consider a compact manifold $M$ with a $(G,X)$-foliation $\mathfrak{F}$, i.e. a geometric structure supported on a foliation $\mathfrak{F}$ of $M$. We call the couple $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. Roughly speaking, the leaves have a $(G,X)$-structure that varies continuously from a leaf to another. Those manifolds arise naturally in many interesting geometric settings. For example, the so-called pp-waves (Example 2.4), and more generally, weakly pp-waves (Example 3.1), are codimension $1$ foliated Lorentzian manifolds with a rigid structure on the leaves. Namely, the leaves of a pp-wave (resp. of a weakly pp-wave) have a special affine structure called “the affine unimodular lightlike geometry” (resp. “the affine lightlike geometry”), introduced in subsections 1.3.1 and 1.3.2 below. Note that $(G,X)$-foliations are briefly mentioned in Thurston’s book [22, p.160], but not precisely defined. Since they arise naturally, it is interesting to consider their completeness. A more precise formulation of Item (B) is the following
Problem 1.8.
Consider a $(G,X)$-structure for which any compact $(G,X)$-manifold is complete, we ask the following question: for a compact $(G,X)$-foliated manifold $M$, are the leaves $(G,X)$-complete? We call this the $(G,X)$-foliated completeness problem.
To the best of our knowledge, the $(G,X)$-foliated completeness problem has not been considered before. Besides the fact that it is itself interesting and natural, there are some particular geometric structures for which the full completeness reduces essentially to a $(G,X)$-foliated completeness problem.
1.3.1. Affine unimodular lightlike geometry
An $(\sf{Aff}(\mathbb{R}^{n}),\mathbb{R}^{n})$-structure on a manifold $M$ is called an affine structure. For a complete survey on affine structures one can see [10, Ch. 6-10]. We will be interested in a rigid version of the full affine geometry, which we call the affine unimodular lightlike geometry. It was first introduced in [14, page 3]. The structure group is a subgroup of the affine group, given by
${\sf{L}}_{\sf{u}}(n,0):=\left\{\begin{pmatrix}1&\alpha^{\top}\\ 0&A\end{pmatrix}\ltimes\mathbb{R}^{n+1}\ |\ \alpha\in\mathbb{R}^{n},\ A\in{\sf{O}}(n)\right\},$
where $\mathsf{O}(n)$ is the orthogonal group of the Euclidean scalar product on $\mathbb{R}^{n}$. The $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry is the maximal affine geometry with two $\mathsf{L}_{\sf{u}}(n,0)$-invariant objects: the vector field $V:=\partial_{x_{1}}$ and a degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+..+dx_{n+1}^{2}$, with kernel $V$. The notation $(n,0)$ refers to the signature of this degenerate Riemannian metric on $\mathbb{R}^{n+1}$, and is useful to distinguish between different signatures in the next sections. However, when the second entry is 00, we simply denote ${\sf{L}}_{\sf{u}}(n)$ instead of $\mathsf{L}_{\sf{u}}(n,0)$.
It turns out that the (flat) connection preserved by ${\sf{L}}_{\sf{u}}(n)$ belongs to what is called in [8] distal connections. It is proved in the same paper that such a connection is complete when the manifold is compact. For the foliated problem (B), we show
Theorem 1.9.
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. The leaves of $\mathfrak{F}$ are $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-complete. In particular, when the codimension is 00, we obtain that a compact $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-manifold is complete.
This gives for this geometry an affirmative answer to $(G,X)$-foliated completeness problem.
One could also consider a “dual” structure to $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where the structure group is defined by
${\sf{L}}_{\sf{u}}^{*}(n):=\left\{\begin{pmatrix}A&\beta\\ 0&1\end{pmatrix}\ltimes\mathbb{R}^{n+1}\ |\ \beta\in\mathbb{R}^{n},\ A\in{\sf{O}}(n)\right\}.$
The linear part of the structure group ${\sf{L}}_{\sf{u}}^{*}(n)$ is the dual of the linear part of ${\sf{L}}_{\sf{u}}(n)$, hence the name of dual geometry. It has an invariant parallel $1$-form (instead of an invariant parallel vector field) and an invariant foliation (instead of only a distribution). For this dual geometry, we establish an analogue statement to Theorem 1.9.
Theorem 1.10.
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}^{*}(n),\mathbb{R}^{n+1})$-foliated manifold. The leaves of $\mathfrak{F}$ are $(\mathsf{L}_{\sf{u}}^{*}(n),\mathbb{R}^{n+1})$-complete.
1.3.2. Affine lightlike geometry
One could relax the $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry by assuming instead of an invariant vector field, only an invariant line field. This geometry also comes from concrete examples known as “weakly pp-waves” (see Example 3.1). In this case, the structure group is
${\sf{L}}(n):=\left\{\begin{pmatrix}\lambda&\alpha^{\top}\\ 0&A\end{pmatrix}\ltimes\mathbb{R}^{n+1}\ |\ \lambda\in\mathbb{R}_{>0},\ \alpha\in\mathbb{R}^{n},\ A\in{\sf{O}}(n)\right\}.$
We call this $({\sf{L}}(n),\mathbb{R}^{n+1})$-geometry the affine lightlike geometry.
This geometry, however, doesn’t satisfy that any compact $({\sf{L}}(n),\mathbb{R}^{n+1})$-manifold is complete. An incomplete example can be obtained simply by taking the product of a Hopf circle and a Euclidean torus. In this case, and using essentially the same techniques as for Theorem 1.9, we characterize the completeness in the following way
Theorem 1.11.
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated compact manifold. The leaves of $\mathfrak{F}$ are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-complete if and only if all geodesics tangent to the $\mathcal{V}$-leaves are affinely (or geodesically) complete , where $\mathcal{V}$ is the $1$-dimensional foliation tangent to the line field induced on any $\mathfrak{F}$-leaf by this structure.
We give an explicit example showing that there are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated compact manifolds, where the leaves are not complete.
1.3.3. Pseudo-Riemannian variants.
A natural generalization of the above situation is the following pseudo-Riemannian variant of the $({\sf{L}}_{\sf{u}}(n,0),\mathbb{R}^{n+1})$-geometry. Namely,
${\sf{L}}_{\sf{u}}(p,q):=\left\{\begin{pmatrix}1&\alpha^{\top}\\ 0&A\end{pmatrix}\ltimes\mathbb{R}^{p+q+1}\ |\ \alpha\in\mathbb{R}^{p+q},\ A\in{\sf{O}}(p,q)\right\},$
where $\mathsf{O}(p,q)$ is the orthogonal group of the pseudo-Riemannian scalar product of signature $(p,q)$. The completeness of a compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-manifold (and its dual geometry) is itself an open question. In a recent paper [3], the author shows that the answer is positive for the $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-geometry, i.e. a compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-manifold is complete. But it turns out that the foliated completeness Problem 1.8 fails for this geometry: we give an example (Example 4.4) of a compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-foliated manifold with incomplete leaves, for any $n\geq 2$.
For the dual geometry we show
Proposition 1.12.
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n-1,1),\mathbb{R}^{n+1})$-manifold. If $\pi_{1}(M)$ is virtually solvable, then $M$ is complete.
From the incomplete Example 4.4, one can in fact obtain compact $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-foliated manifolds as well as compact $({\sf{L}}_{\sf{u}}^{*}(p,q),\mathbb{R}^{p+q+1})$-manifolds with incomplete leaves, when $p,q\geq 1$.
Remark 1.13.
Note that for the compact $({\sf{L}}_{\sf{u}}(n-1,1),\mathbb{R}^{n+1})$-foliated manifolds ($n\geq 2$) with incomplete leaves, constructed in Example 4.4, the structure group preserves a Lorentzian metric. Thus, the question is still open when the holonomy doesn’t preserve a Lorentzian metric. The Latter case naturally arises if we consider “generalized” pp-waves. Namely, they are manifolds with a metric of index $(n,2)$, which have a parallel vector field $V$ with a flat $V^{\perp}$-distribution.
Organization of the paper
The organization of the paper is different from the order presented in the introduction. Section 2 deals with the $(G,X)$-foliated completeness problem in Item (B), in the case of the affine unimodular lightlike geometry and its dual. We prove Theorem 1.9 and Theorem 1.10 which give leafwise completeness results. Then we study the case of the affine lightlike geometry in Section 3: we prove Theorem 1.11 that characterizes the leafwise completeness, and give leafwise-incomplete examples. Section 4 deals with the pseudo-Riemannian variants of the previous structures. In Section 5, we deal with the foliated completeness problem in Item (A), and prove Theorem 5.1. Finally, Section 6 is devoted to the study of the global completeness question in the presence of a null Killing field, focusing on dimension $3$ in Section 7.
Acknowledgement
We would like to warmly thank Abdelghani Zeghib and Souheib Allout for the helpful discussions, that made this work possible. The first author is also grateful to Ines Kath for her support and encouragements, and for the helpful remarks on the early drafts of this paper. The second author would like to warmly thank Miguel Sanchez for his interest in her work, and for pointing out to her that there is no known incomplete compact Lorentzian manifold with a null Killing field, which motivated this part of the work.
The second author is partially supported by the grants, PID2020-116126GB-I00
(MCIN/ AEI/10.13039/501100011033), and the framework IMAG/ Maria de Maeztu,
CEX2020-001105-MCIN/ AEI/ 10.13039/501100011033.
2. $(G,X)$-foliated completeness problem: affine unimodular lightlike geometry
2.1. $(G,X)$-structures.
Let $G$ be a Lie group, and $X$ a space on which $G$ acts smoothly transitively, then, one has $X=G/H$, where $H$ is a closed subgroup of $G$. A $(G,X)$-structure on $M$ is an atlas on $M$, where the chart transitions are restrictions of elements of $G$. A manifold with a $(G,X)$-structure is called a $(G,X)$-manifold. A $(G,X)$-manifold $M$ gives rise to a local diffeomorphism $D$, called the developing map, from the universal cover to the model space $X$, together with a representation $\rho:\pi_{1}(M)\to G$, called the holonomy representation (see [10, p. 100] or [23, p. 34]). Moreover, the developing map and the holonomy representation satisfy the following equivariance property
$D(\gamma x)=\rho(\gamma)D(x),$
for any $x\in\widetilde{M}$ and $\gamma\in\pi_{1}(M)$.
Definition 2.1 (Completeness in the sense of $(G,X)$-structures).
Let $M$ be a $(G,X)$-manifold, we say that $M$ is complete if the developing map $D:\widetilde{M}\to X$ is a covering map.
Fact 2.2.
Let $M$ be $(G,X)$-manifold. Assume that $X$ is complete with respect to a $G$-invariant connection. Then $M$ is geodesically complete (with respect to the “pullback” connection) if and only if $M$ is $(G,X)$-complete.
2.2. $(G,X)$-foliation
We introduce here the notion of a $(G,X)$-foliation. Examples are given in the next sections. A smooth manifold $M$ with a foliation $\mathfrak{F}$ will be called a foliated manifold and denoted by a couple $(M,\mathfrak{F})$.
Definition 2.3 ($(G,X)$-foliation).
We say that a manifold $M$ has a $(G,X)$-foliation if $M$ is a foliated manifold such that the leaves of $\mathfrak{F}$ are modeled on the $(G,X)$-geometry. More precisely, for every $p\in M$ there exist an open neighborhood $U$ and a diffeomorphism (into the image) $\varphi:U\to I\times X$ such that every leaf the foliation of $U$ (induced by $\mathfrak{F}$) is mapped to a level $(t,V_{t})\subset I\times X$. And if $(\varphi_{i},U_{i})$ and $(\varphi_{j},U_{j})$ are two such maps, the transition map $\varphi_{ij}:=\varphi_{j}\varphi_{i}^{-1}$ on each level $(t,V_{t})\subset\varphi_{i}(U_{i}\cap U_{j})$ is the restriction of an element $g_{t}\in G$. Moreover, the $g_{t}$ (necessarily unique) varies continuously from a level to another.
Pullback of $G$-invariant tensor fields. Consider $(M,\mathfrak{F})$ a $(G,X)$-foliated manifold. Let $T$ be a tensor field on $X$ which is $G$-invariant. One can associate a tensor field $T_{M}$ on $M$ in the following way. For any point $p\in M$, we take a chart around $p$ and we pullback the tensor field $T$ in a neighborhood of $p$. We claim that we get a well defined tensor field $T_{M}$ on $M$. Indeed, the transition maps, when restricted to a leaf, are restrictions of an element of $G$, and $T$ is $G$-invariant. Moreover, by construction, $T_{M}$ is continuous on $M$. We will refer to $T_{M}$ as the “pullback” of $T$.
2.3. Affine unimodular lightlike geometry.
Recall that the affine unimodular lightlike geometry is the one modeled on $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$, where $\mathsf{L}_{\sf{u}}(n)$ is the subgroup of $\sf{Aff}(\mathbb{R}^{n+1})$ preserving the degenerate Riemannian metric $h:=dx_{2}^{2}+dx_{3}^{2}+..+dx_{n+1}^{2}$, and the isotropic vector field $V:=\partial_{x_{1}}$. In this subsection, $V$ and $h$ will always refer to these ${\sf{L}}_{\sf{u}}(k)$-invariant objects. The foliation by the leaves of $V$ will be denoted by $\mathcal{V}$.
Example 2.4 (pp-waves are $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifolds).
A pp-wave is a Lorentzian manifold $(M,V)$ with a parallel null vector field $V$, such that the (integrable) parallel distribution $V^{\perp}$ is flat with respect to the induced connection (this distribution defines a foliation $\mathfrak{F}$ which is totally geodesic). The leaves of $\mathfrak{F}$ are endowed naturally with a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-structure, which varies smoothly. Indeed, it is well known that the data of a flat affine connection on a smooth manifold is equivalent to have an $(\sf{Aff}(\mathbb{R}^{n+1}),\mathbb{R}^{n+1})$-structure on it. Hence, pp-waves have an induced affine structure on the leaves of $\mathfrak{F}$. However, due to the existence of a parallel vector field $V$ and the fact that $V^{\perp}$ has a natural degenerate Riemannian metric with kernel $\mathbb{R}V$, the leaves have in fact a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry. So pp-waves give natural examples of $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifolds.
2.4. The $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated completeness
This subsection is devoted to the study of the foliated completeness problem, we show the following
Theorem 2.5.
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated compact manifold. Then the leaves of $\mathfrak{F}$ are $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-complete.
The proof, presented in Paragraph 2.4.2, is based on the theory of Riemannian submersions, which was developed by B. O’Neil in [19] and A. Grey in [13]. We recall the main results we will need in the next subsection.
2.4.1. Prerequisites on Riemannian submersions
Let $(E,g)$ and $(B,g^{\prime})$ be smooth Riemannian manifolds, where $\dim E\geq\dim B$. Let $\pi:E\to B$ be a smooth submersion, the regular value theorem implies that the fiber $\pi^{-1}(x)$ of any point $x\in B$ is a closed smooth submanifold of $E$ of dimension $\dim E-\dim B$. Let $p\in E$, we denote by $\mathcal{K}_{p}$ the tangent space at $p$ of the fiber of $\pi$, and $\mathcal{H}_{p}$ its orthogonal complement at $p$ with respect to the metric $g$. Thus, we can decompose the tangent space at $p$ as $T_{p}E=\mathcal{K}_{p}\oplus\mathcal{H}_{p}$. The distribution $\mathcal{K}$ is called the vertical distribution, it is integrable, and the leaves are the fibers of $\pi$. The distribution $\mathcal{H}$ is called the horizontal distribution; generically, it is far from being integrable.
Definition 2.6.
Let $\pi:(E,g)\to(B,g^{\prime})$ be a smooth submersion, we say that $\pi$ is a Riemannian submersion if the differential of $\pi$ induces an isometry from the horizontal distribution to the tangent space of $B$. In other words, $d_{p}\pi$ is an isometry from $\mathcal{H}_{p}$ to $T_{\pi(p)}B$.
The horizontal distribution of a Riemannian submersion.
In all this paragraph, we consider $\pi:(E,g)\to(B,g^{\prime})$ a Riemannian submersion, and $\mathcal{H}$ the corresponding horizontal distribution. A vector $v\in T_{p}M$ is said to be horizontal if $v\in\mathcal{H}_{p}$. And a curve $\gamma:I\to E$ is said to be horizontal if for all $t\in\mathbb{R}$, $\gamma^{\prime}(t)$ is a horizontal vector, i.e. $\gamma^{\prime}(t)\in\mathcal{H}_{\gamma(t)}$.
Fact 2.7 (O’Neil, [20] Proposition 1.10).
Let $\gamma:I\to E$ be a geodesic. If the tangent vector $\gamma^{\prime}(t_{o})$ is horizontal, then $\gamma$ is horizontal. In other words, the horizontal distribution is totally geodesic.
Proposition 2.8 (Corollary 1.1 page 26, [20]).
Let $\pi:(E,g)\to(B,g^{\prime})$ be a Riemannian submersion. Consider $\gamma:I\subset\mathbb{R}\to E$ a parameterized horizontal curve with constant speed, i.e. for all $t\in I$ we have $\gamma^{\prime}(t)\in\mathcal{H}_{\gamma(t)}$ and $g_{\gamma(t)}(\gamma^{\prime}(t),\gamma^{\prime}(t))=1$. Then, $\gamma$ is a geodesic in $E$ if and only if $\pi\circ\gamma$ is a geodesic in $B$.
Fact 2.9 (Proposition 3.2 [15]).
Let $\pi:(E,g)\to(B,g^{\prime})$ be a Riemannian submersion. If the total space $E$ is complete, then $E$ is a fiber bundle over $B$.
2.4.2. Proof of Theorem 2.5
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. For simplicity, we will assume that the leaves are of codimension $1$, i.e. $\dim M=n+2$ (although the proof works the same way for higher codimension). Recall that on $\mathbb{R}^{n+1}$ we have two $\mathsf{L}_{\sf{u}}(n)$-invariant objects: a vector field $V$ and a degenerate Riemannian metric $h$. Consider $V_{M}$ the pullback of $V$, it is a continuous non-vanishing vector field on $M$ tangent to $\mathfrak{F}$. Consider also $h_{M}$ the degenerate metric given by the pullback of $h$ on the subbundle $T\mathfrak{F}$. Denote by $\mathcal{V}_{M}$ the foliation tangent to $V_{M}$.
Choose an arbitrary Riemannian metric $g_{o}$ on $M$, and decompose the tangent space $TM$ orthogonaly with respect to $g_{o}$ as follows,
$TM=T\mathfrak{F}\oplus\mathbb{R}=\mathbb{R}V_{M}\oplus\mathcal{S}\oplus\mathbb{R}Z,$
where $Z$ is a unit vector field with respect to $g_{o}$ (it exists, up to taking a double cover).
In order to prove Theorem 2.5, we need to show (using Fact 2.2) that the developing map $D$ is a diffeomorphism from the universal cover of any leaf of $\mathfrak{F}$ onto the affine space $\mathbb{R}^{n+1}$. This will be done by constructing a particular Riemannian submersion, then we apply Subsection 2.4.1, especially, Fact 2.9.
Constructing the Riemannian submersion.
A first step toward the construction of the Riemannian submersion is to define a Riemannian metric $g$ on $M$ compatible with $h_{M}$ in the following sense
$g(V_{M},V_{M})=g(Z,Z)=1,\ g(V_{M},Z)=0,$
$g(V_{M},X)=0,\ g(Z,X)=0,\ g(X,Y)=h_{M}(X,Y)$
for all $X,Y\in\mathcal{S}$. The metric $g$ is indeed Riemannian since the restriction of $h_{M}$ to the distribution $\mathcal{S}$ is non-degenerate and Riemannian (remember that $\mathcal{S}$ is transversal to the kernel of $h_{M}$, namely, the direction of $V_{M}$).
Fact 2.10.
Let $(M,\mathfrak{F},g)$ be a compact foliated Riemannian manifold. Then any leaf $\mathcal{F}$ with the induced Riemannian metric is geodesically complete.
Corollary 2.11.
Let $\mathcal{F}$ be a leaf of $\mathfrak{F}$. Then $(\mathcal{F},g_{|_{\mathcal{F}}})$ is geodesically complete.
Let (~$(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}_{p}}}})$ be the universal cover of $\mathcal{F}$ together with the lifted metric. The Riemannian manifold (~$(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}})$ is geodesically complete, so we have a developing map $D:\widetilde{\mathcal{F}}\to\mathbb{R}^{n+1}$ which is a local diffeomorphism. Denote by $\mathfrak{p}$ the projection from $\mathbb{R}^{n+1}$ to the space of $\mathcal{V}$-leaves, i.e. $\mathfrak{p}:\mathbb{R}^{n+1}\to\mathbb{R}^{n+1}/\mathcal{V}\cong\mathbb{R}^{n}$. The space of $\mathcal{V}$-leaves is $\mathbb{R}^{n}$ equipped with the flat Riemannian metric, given by the projection of $h$. We denote the projected metric by $\bar{h}$. Define $\pi:=\mathfrak{p}\circ D$, this is a smooth submersion from $\widetilde{\mathcal{F}}$ to $B:=\pi(\widetilde{\mathcal{F}})$, which is an open subset of $\mathbb{R}^{n}$.
Lemma 2.12.
The map $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(B,\bar{h}_{|_{B}})$ is a Riemannian submersion.
Proof.
The horizontal distribution on $\widetilde{\mathcal{F}}$ is given by $\widetilde{S}$ (the lift of $S$). Let $x\in\widetilde{\mathcal{F}}$ and $y:=D(x)$, then $E_{y}:=d_{x}D(\widetilde{S}_{x})$ is transversal to the kernel of $h_{y}$ (i.e. to $\mathbb{R}V$). By construction, the metric induced on $\widetilde{S}_{x}$ by $g_{x}$ is isometric to the metric induced on $E_{y}$ by $h_{y}$. The lemma follows then from the fact that the metric induced by $h_{y}$ on any transversal is isometric to $(T_{y}\mathbb{R}^{n+1}/\mathbb{R}V,\bar{h})$. ∎
Corollary 2.13.
We have $B=\mathbb{R}^{n}$.
Proof.
Let $\gamma:I\to\widetilde{\mathcal{F}}$ be a horizontal geodesic (recall that the horizontal distribution is totally geodesic by Fact 2.7). Then $\gamma$ is complete by Corollary 2.11. By Proposition 2.8, the projection of $\gamma$ by $\pi$ is a geodesic in the base space, which is an open subset of the flat euclidean space, hence $\pi\circ\gamma$ is a line segment. Since $\gamma$ is complete, this line segment is a complete affine line. This is true for any horizontal geodesic in all possible directions, hence $B=\mathbb{R}^{n}$. ∎
Corollary 2.14.
The map $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(\mathbb{R}^{n},\bar{h})$ is a trivial topological fiber bundle, i.e. we have ~$\widetilde{\mathcal{F}}\cong\mathbb{R}^{n}\times F$, with $F=\pi^{-1}(0)$.
Proof.
By Fact 2.9, we have that $\pi:(\widetilde{\mathcal{F}},\widetilde{g_{|_{{\mathcal{F}}}}}))\to(\mathbb{R}^{d},\bar{h})$ is a topological fiber bundle. Since the base space is contractible, the bundle is trivial (a product bundle). ∎
Theorem 2.15 (Theorem 2.5).
Let $(M,\mathfrak{F})$ be a $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated compact manifold. Then the leaves of $\mathfrak{F}$ are $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-complete.
Proof.
Let $\mathcal{F}$ be a leaf of $\mathfrak{F}$, with $D:\widetilde{\mathcal{F}}\to\mathbb{R}^{n+1}$ a developing map. We will show that $D$ is a diffeomorphism. We start by the surjectivity. Recall that we have a (surjective) Riemannian submersion $\pi:\widetilde{\mathcal{F}}\to\mathbb{R}^{n}$, where $\mathbb{R}^{n}$ is the Euclidean space. Let $z\in\mathbb{R}^{n+1}$ be any point, let $x=\pi(x_{o})$ for some $x_{o}\in\widetilde{\mathcal{F}}\smallsetminus\pi^{-1}(\{z\})$. Now, consider the line segment $\gamma:=[x,\mathfrak{p}(z)]$ in $\mathbb{R}^{n}$, we set $z_{o}$ to be the other endpoint of the horizontal lift $\tilde{\gamma}$ starting at $x_{o}$. Since $\pi(z_{o})=\mathfrak{p}(z)$ we see that $D(z_{o})\in\mathcal{V}_{z}$. Moreover, the developing map sends the $\widetilde{\mathcal{V}}_{M}$ leaf containing $z_{o}$ to the $\mathcal{V}$ leaf containing $z$, i.e. $D(\widetilde{\mathcal{V}}_{M,z_{o}})\subset\mathcal{V}_{z}$. We claim that the restriction of $D$ to any $\widetilde{\mathcal{V}}_{M}$ is injective. Indeed, since the vector field $\widetilde{V}_{M}$ is the pullback of $V$ via $D$, we have that $D$ is equivariant with respect to their flows, i.e. for $\tilde{p}\in\widetilde{F}$ we have
$D(\phi_{\widetilde{V}_{M}}^{t}(\tilde{p}))=\phi_{V}^{t}(D(\tilde{p})).$
Moreover, the flow of $V$ acts freely on the affine space, in particular on its orbits (affine lines). The claim follows. Now, since $M$ is compact the flow of $\widetilde{V}_{M}$ is complete, so, the restriction of developing map to a $\widetilde{\mathcal{V}}_{M}$ leaf is also surjective. Hence, there exists some $z^{\prime}\in\widetilde{\mathcal{V}}_{M,z_{o}}$ such that $D(z^{\prime})=z$.
For the injectivity of $D$, assume that there are two distinct points $x_{o},y_{o}\in\widetilde{\mathcal{F}}$ (necessarily in different $\widetilde{\mathcal{V}}_{M}$ leaves) such that $D(x_{o})=D(y_{o})$. Hence, $\pi(x_{o})=\pi(y_{o})=x$, which means that the fiber of $x$ contains at least two $\widetilde{\mathcal{V}}_{M}$ leaves. On other hand, Corollary 2.14 together with the fact that $\widetilde{\mathcal{F}}$ is connected implies that the fiber is necessarily connected. Therefore contradiction. ∎
Corollary 2.16.
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. Then the leaves of $\mathfrak{F}$ are geodesically complete with respect to the flat affine connection.
Proof.
On a leaf $\mathcal{F}$ of $\mathfrak{F}$ we can define the “pullback” of the flat affine connection (since it is invariant by $\mathsf{L}_{\sf{u}}(n)$). However, since $\mathcal{F}$ is $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-complete we deduce by Fact 2.2 that it is geodesically complete. ∎
Corollary 2.17 (The dual geometry).
Let $(M,\mathfrak{F})$ be a compact $(\mathsf{L}_{\sf{u}}^{*}(n),\mathbb{R}^{n+1})$-foliated manifold. Then the leaves of $\mathfrak{F}$ are $(\mathsf{L}_{\sf{u}}^{*}(n),\mathbb{R}^{n+1})$-complete.
Proof.
Recall that the invariant objects for this structure are the $1$-form $\omega:=dx_{n+1}$ and the codimension $1$ foliation $\mathfrak{G}$ tangent to the kernel of this form. Denote by $\mathfrak{G}_{M}$ the foliation on $M$, tangent to $\mathfrak{F}$, given by the pullback of $\mathfrak{G}$. And denote by $\omega_{M}$ the $1$-form on $M$ given by the pullback of $\omega$. Similarly to the proof of Theorem 2.5, we show that the leaves of $\mathfrak{G}_{M}$ are complete. On the other hand, there is a vector field $W$ of $M$ such that $\omega_{M}(W)=1$. Indeed, $W$ can be chosen as a global section of a tangent line bundle, this global section exists since the fibers are contractible. Consider the $1$-from $\widetilde{\omega}_{M}$ and the vector field $\widetilde{W}$ on $\widetilde{M}$. Let $\tilde{p}\in\widetilde{M}$. Since $\omega$ is $\mathsf{L}_{\sf{u}}^{*}(n)$-invariant, we have the following formula
$D\circ\phi_{\widetilde{W}}^{t}(\widetilde{\mathcal{G}}_{M}(\tilde{p}))=\phi_{\partial x_{n+1}}^{t}\circ D(\widetilde{\mathcal{G}}_{M}(\tilde{p}))=\phi_{\partial x_{n+1}}^{t}(\mathcal{G}(D(\tilde{p}))),$
where $\phi_{\widetilde{W}}^{t}$ is the flow of $\widetilde{W}$, where $\mathcal{G}(p)$ (resp. $\mathcal{G}_{M}(p)$) is the leaf of $\mathfrak{G}$ (resp. $\mathfrak{G}_{M}$) at $p$. From this we see that $D$ is surjective since $\widetilde{W}$ is complete (by compactness of $M$), and injective since the $\widetilde{W}$-flow acts transitively on the space of $\widetilde{\mathfrak{G}}_{M}$-leaves. ∎
2.5. Application to compact pp-waves
The so-called pp-waves are a special class of Lorentz- ian manifolds with a parallel null vector field $V$, with a great importance in General Relativity. It was proved in [17] that a compact pp-wave is complete. The completeness of the leaves of the codimension $1$ totally geodesic foliation $\mathfrak{F}$ tangent to $V^{\perp}$ was a crucial step in the proof. On the other hand, it is observed in Example 2.4 that the leaves of $\mathfrak{F}$ have a natural $(\mathsf{L}_{\sf{u}}(k),\mathbb{R}^{k+1})$-structure. We then recover, applying Corollary 2.16, the completeness result along the leaves of the codimension $1$ foliation $\mathfrak{F}$. We see that the completeness of the $\mathfrak{F}$-leaves of a pp-wave is independent of the presence of a Lorentzian ambient metric, and is therefore purely a leafwise completeness problem!
3. $(G,X)$-foliated completeness problem: affine lightlike geometry
We relax the geometric structure $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$ by assuming that only the direction of the vector field $V:=\partial_{x_{1}}$ is preserved by the structure group. In other words, the data is an invariant line field rather than an invariant vector field. The geometry in this case is the affine lightlike geometry $(\mathsf{L}(n),\mathbb{R}^{n+1})$. We denote again by $\mathcal{V}_{M}$ the foliation tangent to the line field induced on $M$ by this structure.
Example 3.1 (weakly pp-waves are $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated manifolds).
Let $(M,l)$ be a “weakly pp-wave” of dimension $n+2$, that is, a Lorentzian manifold with a null parallel line field $l$, for which the (totally geodesic) foliation $\mathfrak{F}$ orthogonal to $l$ is flat with respect to the induced connection. Then, the leaves of $\mathfrak{F}$ have a natural affine lightlike geometry. So weakly pp-waves give natural examples of $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated manifolds.
We ask whether a completeness statement analogous to Theorem 2.5 still holds for this geometric structure $({\sf{L}}(n),\mathbb{R}^{n+1})$. Notice that the methods from the previous case of $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry give straightforwardly that the developing map $D$ (on the universal cover of some leaf) is injective, and surjective on $\mathcal{F}/\mathcal{V}_{M}$, but fail in the surjectivity along the $\mathcal{V}_{M}$-leaves. This failure is due to lack of an invariant vector field. Indeed, the key point for surjectivity along the $\mathcal{V}_{M}$-leaves was that the $V_{M}$-flow is equivariant and complete. Formulating this observation we get the following
Theorem 3.2.
Let $(M,\mathfrak{F})$ be a compact $({\sf{L}}(n),\mathbb{R}^{n+1})$-foliated manifold. A leaf of $\mathfrak{F}$ is complete if and only if the geodesics tangent to $\mathcal{V}_{M}$ are complete.
However, it turns out that the incomplete situation can occur. In the following we give some examples illustrating this. We present two kind of examples: the first kind is an affine manifold with a codimension $1$ foliation that supports an $({\sf{L}}(n),\mathbb{R}^{n+1})$-structure, and the second kind is a “weakly pp-wave”.
Example 3.3 ($\mathsf{Hopf^{\lambda}_{1}}\times\mathbb{T}^{2}$).
Let $\lambda>0$. Consider the open subset of $\mathbb{R}^{3}$ given by $U:=\{(v,x,u),v>0\}$, and take the quotient of $U$ by $\Gamma$ generated by
$\gamma_{1}$ $:(v,x,u)\to(\lambda v,x,u)$
$\gamma_{2}$ $:(v,x,u)\to(v,x+1,u)$
$\gamma_{3}$ $:(v,x,u)\to(v,x,u+1).$
Here, $\Gamma$ is a subgroup of $\sf{Aff}(\mathbb{R}^{3})$ acting properly discontinuously and freely on $U$, with a compact quotient. So $\Gamma\backslash U$ is a compact affine manifold: the product of a Euclidean torus and an Hopf circle of affine length $\lambda$. Now, equip $\mathbb{R}^{3}$ with the Lorentzian flat metric $2dudv+dx^{2}$. The subgroup $\Gamma$ does not preserve the flat metric, but it preserves the null direction $\mathbb{R}V$, where $V=\partial_{v}$, together with the degenerate Riemannian metric $dx^{2}$ on the leaves of $V^{\perp}$. Hence the foliation tangent to $V^{\perp}$ is a $({\sf{L}}(1),\mathbb{R}^{2})$-foliation on $\Gamma\backslash U$. Note that if $\Gamma$ is a subgroup of the Poincaré group, i.e. preserves the Lorentzian flat metric, then $U$ is necessarily $\mathbb{R}^{3}$ and the quotient is complete due to Carriére’s result. We see that in this case, having a Lorentzian ambient structure is not artificial regarding the foliated-completeness question.
Example 3.4 (Clifton-Pohl torus).
Let $(T_{CP},g):=(\mathbb{R}^{2}\smallsetminus\{(0,0)\},\frac{2dxdy}{x^{2}+y^{2}})/(x,y)\sim 2(x,y)$ denote the Clifton-Pohl torus. Let $\mathcal{N}$ be one of its two null foliations. Then the tangent bundle of $\mathcal{N}$ defines a parallel null line on $T_{CP}$. It is known that the leaves of the foliation are incomplete.
Example 3.5 (Higher dimension).
(1) Consider the product $M=T_{CP}\times\mathbb{S}^{1}$, with Lorentzian metric $g+dt^{2}$. Denote by $\mathcal{N}$ the null foliation on $M$ coming from one of the null foliations of $T_{CP}$. Its tangent bundle defines a null parallel line field on $M$. The orthogonal leaves $\mathcal{F}=\mathcal{N}^{\perp}$ are given by $\mathcal{N}\times\mathbb{S}^{1}$. They are flat (with respect to the induced connection) and incomplete. In this example, a leaf of $\mathfrak{F}$ is either a cylinder or a torus, and the universal cover of a leaf develops onto ×$\mathbb{R}\times I$, where $I$ is some open interval in $\mathbb{R}$ which is either affinely equivalent to ]0,1[]0,1[] 0 , 1 [ or to ]0,+[]0,$\infty$[] 0 , + ∞ [ (both cases occur).
(2) One can deform the previous example in order to have simply connected $\mathfrak{F}$-leaves. Consider $X=\mathbb{R}^{3}\smallsetminus\{(0,0\}\times\mathbb{R}$ with Lorentzian metric $g+dt^{2}$. Its isometry group is $G=\mathbb{R}\times\mathbb{R}$ (up to finite index), where the first $\mathbb{R}$-factor acts by a $1$-parameter group of homotheties on the Clifton-Pohl metric of $\mathbb{R}^{2}\smallsetminus\{(0,0)\}$, and the second $\mathbb{R}$-factor acts by translation on $\mathbb{R}$. The group $G$ preserves the foliation $\mathcal{N}\times\mathbb{R}$ of $X$. Consider a discrete subgroup of $G$ acting properly cocompactly on $X$, and not intersecting the second $\mathbb{R}$-factor. The quotient is a Lorentzian torus with a parallel null line field, and having simply connected orthogonal leaves.
[…]
In the above examples, incompleteness occurs along geodesics tangent to $\mathcal{N}$, but also on spacelike geodesics not tangent to $\mathbb{S}^{1}$ (a flat band of the form ×$\mathbb{R}\times]0,1[\subset\mathbb{R}^{2}$ contains spacelike lines that leave the band). However, in this example, any spacelike geodesic tangent to a leaf is contained in an infinite band (infinite here is for the Riemannian transverse structure).
4. Foliated $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-geometry and its dual
In this section we show that the foliated completeness Problem 1.8 is not true for the $({\sf{L}}_{\sf{u}}(1,k),\mathbb{R}^{k+2})$-structure and its dual, by constructing concrete incomplete examples.
4.1. Compact $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-manifolds
In a recent paper [3], it is shown that a compact $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-manifold is complete. The proof uses techniques from [7, 9, 6].
4.2. Compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifolds
The dual structures were mentioned earlier by Carrière in [6, Theorem 3.2.1]. Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. This structure on $M$ gives rise naturally to a codimension $1$ totally geodesic Lorentzian foliation with flat leaves. In [6, Theorem 3.2.1], it was shown that the developing map on those leaves is injective. Here we use the extra invariant $1$-form to show that the developing map of $M$ is itself injective.
Fact 4.1.
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. The developing map $d$ is injective. As a result, the holonomy representation is injective, and the fundamental group is isomorphic to a discrete subgroup of ${\sf{L}}_{\sf{u}}^{*}(k,1)$.
Proof.
The proof for the injectivity of $d$ is the same as in Corollary 2.17. The rest is an immediate consequence. ∎
This allows to show
Corollary 4.2 (Solvable implies complete).
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. If $\pi_{1}(M)$ is virtually solvable, then $M$ is complete.
Proof.
First, we remark that $M$ is aspherical. Indeed, $\widetilde{M}$ fibers over $\mathbb{R}$ with a contractible fiber (corresponding to the Lorentzian flat leaves, see [6, Proposition 3.2.2]). Moreover, since $\pi_{1}(M)$ is virtually isomorphic to a discrete subgroup of a connected Lie group, $\pi_{1}(M)$ is virtually solvable if and only if it is virtually polycyclic [12, Section 1]. Hence, if $\pi_{1}(M)$ is virtually solvable we have $\mathsf{Rank}(\pi_{1})=\mathsf{cd}(\pi_{1})$=$\dim(M)$. However, $M$ has a parallel volume. The claim follows using [11, Theorem 3.5]. ∎
Corollary 4.3.
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{1+2})$-manifold. Then $M$ is complete.
Proof.
Fact 4.1 implies that the holonomy is injective. Hence, the $\pi_{1}(M)$ is isomorphic to a subgroup of ${\sf{L}}_{\sf{u}}^{*}(1,1)$ Clearly, ${\sf{L}}_{\sf{u}}^{*}(1,1)$ is solvable. So, the claim follows from Corollary 4.2. ∎
4.3. The foliated problem
Here we construct an example (Example 4.5) where Problem 1.8 fails. The key fact that makes the problem to fail is the existence of compact manifolds with an incomplete foliated flat Lorentzian structure (Example 3.3).
Example 4.4 (Compact Lorentz manifold with incomplete flat Lorentzian leaves).
Consider ${\mathsf{Sol}}=\mathbb{R}\ltimes\mathbb{R}^{2}$, with its Lie brackets $[A,X]=X$, $[A,Y]=-Y$. Endow it with the left invariant Lorentzian metric given in the bases $(A,X,Y)$ by
$q=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{pmatrix}.$
This is in fact a (non-flat) plane wave metric on ${\mathsf{Sol}}$ (normal-(((see [4, Theorem 1.4])normal-))). The left invariant distribution induced by ${\sf{span}}(A,X)$ is integrable, and a simple computation shows that it is flat, so that the induced metric is a left invariant Lorentzian flat metric on $\sf{Aff}(\mathbb{R})$, which is then incomplete. In fact, the above metric on ${\mathsf{Sol}}$ is given explicitly in coordinates in [4, Theorem 1.4], and the incomplete flat foliation there corresponds to the levels of the $x$-coordinate. Now, consider the left quotient of ${\mathsf{Sol}}$ by some lattice $\Gamma$. The left action of $\Gamma$ preserves this foliation, which is then well defined in the compact quotient $\Gamma\backslash{\mathsf{Sol}}$. Its leaves are incomplete, and develop into the half Minkowski plane. This gives an example of an incomplete foliated $(G,X)$-structure, where the geometry here is the flat Lorentzian structure. It is known that this structure is complete when the manifold is compact (Carrière theorem). This example shows in particular that a foliated Carrière theorem is not true.
Example 4.5 (Incomplete foliated $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-structure).
Consider the product ${\mathsf{Sol}}\times\mathbb{R}$, with the basis $(A,X,Y,V)$ of the Lie algebra, where $V$ is the basis for the $\mathbb{R}$-factor. Define a left invariant Lorentzian metric on it such that: ${\mathsf{Sol}}$ is endowed with the metric given in the previous example, $\mathbb{R}$ is Riemannian, and the two spaces are orthogonal. Namely, in the basis $(A,X,Y,V)$
$q=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}.$
The leaves of the foliation tangent to the left invariant distribution induced by ${\sf{span}}(A,X,V)$ are flat Lorentzian, and they admit a parallel spacelike vector field $V$. Hence, the leaves are modeled on the $(G,X)$-structure, where the group structure in the basis $(V,A,X)$ is given by
$G=\left\{\begin{pmatrix}1&0&0\\ 0&e^{t}&0\\ 0&0&e^{-t}\end{pmatrix}\ltimes\mathbb{R}^{3}\ |\ t\in\mathbb{R}\right\}.$
This is a subgeometry (stiffened) of the $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-geometry. The previous example shows that the leaves are incomplete. This gives an example of an incomplete foliated $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-structure. This example generalizes to a ${\sf{L}}_{\sf{u}}(n-1,1)$-foliated structure.
Remark 4.6 (Incomplete foliated $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$-structure).
The example above shows that Problem 1.8 also fails for the dual $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$-geometry. Indeed, up to permute the basis, the structure group becomes
$\left\{\begin{pmatrix}e^{t}&0&0\\ 0&e^{-t}&0\\ 0&0&1\end{pmatrix}\ltimes\mathbb{R}^{3}\ |\ t\in\mathbb{R}\right\},$
which is also a subgeometry of $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$.
5. Foliated lightlike geometry: leafwise completeness
5.1. Case of a totally geodesic lightlike foliation of a Lorentzian manifold
Let $(M,\mathfrak{F})$ be a compact Lorentzian manifold with a totally geodesic lightlike foliation $\mathfrak{F}$. Each leaf has a degenerate Riemannian metric with $1$-dimensional kernel: we denote by $\mathcal{N}$ the foliation tangent to the kernel.
Theorem 5.1.
Let $(M,\mathfrak{F})$ be a compact Lorentzian manifold with a codimension $1$ totally geodesic lightlike foliation $\mathfrak{F}$. Then the leaves of $\mathfrak{F}$ are complete, with respect to the induced connection, if and only if the (null) geodesics tangent to $\mathcal{N}$ are complete.
Geodesic equation
Near any point $p\in M$, there exists local coordinates $(v,u,x=(x_{1},...,x_{n}))$ where the metric has the following form (see [5, Proposition 2.2]):
$g=2dudv+H(v,u,x)du^{2}+\sum_{i=1}^{n}W_{i}(v,u,x)dudx^{i}+\sum_{i,j}h_{ij}(u,x)dx^{i}dx^{j}.$
Let $V:=\partial_{v}$. In those coordinates, we have $\nabla_{V}V=0$, so the flow of $V$ parameterizes the null geodesics in $\mathcal{N}$. The leaves of $\mathfrak{F}$ are given by the levels of $u$.
For a geodesic tangent to $\mathfrak{F}$, the $u$-coordinate is constant along the geodesic, so that $\dot{u}=0$. For such a geodesic, the geodesic equation simplifies into the following
$\ddot{x}^{k}(t)+\Gamma_{ij}^{k}(u_{0},x)\dot{x}^{i}\dot{x}^{j}=0$
$\ddot{v}(t)+\Gamma_{ij}^{v}(u_{0},x)\dot{x}^{i}\dot{x}^{j}=0$
$u=u_{0}$
We see that the coefficients in the geodesic equations are time independent.
Let $E$ be a smooth $n$-dimensional distribution on $M$, such that for any $p\in M$, $E_{p}$ is a non-degenerate Riemannian subspace of $T_{p}M$ tangent to $\mathfrak{F}$. For every $p\in M$, there exists an open neighborhood $O_{p}$ of 00 in $E_{p}$ such that the exponential map $\exp_{p}:E_{p}\subset T_{p}M\to M$ restricted to $E_{p}$ is a diffeomorphism from $O_{p}$ onto its image in $M$. Set $S_{p}:=\exp_{p}O_{p}$; it is an $n$-submanifold through $p$. For every $p\in M$, consider a normal coordinate system $(x^{1},...,x^{n})$ on $S_{p}$. If $\rho=\sum(x^{i})^{2}$, then for $\delta>0$ sufficiently small, $B_{p}(\delta)=\{q\in S_{p},\rho(q)<\delta\}$ is a normal neighborhood of $p$ in $S_{p}$ diffeomorphic to an open ball $B_{n}(0,\delta)\subset\mathbb{R}^{n}$.
Since $M$ is compact, a continuity argument shows that there exists $r>0$ such that for any $p\in M$, we can take $B_{p}$ above to be the open ball of radius $r$ and center $p$ in $S_{p}$. Now, by compactness of $M$ and the fact that the null geodesics in the kernel of $\mathfrak{F}$ are complete, there exists a positive constant, which we denote again by $r$, such that every point $p\in M$ admits a coordinate chart $(u,v,x^{1},...,x^{n}):U_{p}\subset M\to I\times J\times B_{n}(0,r)$ defined on an open neighborhood $U_{p}$ of $p$, such that $|I|=|J|=r$. Henceforth, to each point $p\in M$ we associate a coordinate chart as before. We can assume without loss of generality that $r=1$. Denote by $h_{u}$ the Riemannian metric induced on the slice $\{u\}\times\{v\}\times B_{n}(0,r)$. It does not depend on $v$.
Remark 5.2.
Let $\gamma$ be a geodesic tangent to $\mathfrak{F}$, such that $\gamma(0)=p\in M$. Let $(u,v,x)$ be a system of coordinates around $p$ as above. Then the $h_{u}$-norm $||\dot{x}(t)||_{u}$ is independent of the coordinate system, since it coincides with $g(\dot{\gamma}(t),\dot{\gamma}(t))$. Moreover, the second order differential equation on $x(t)$ is exactly the geodesic equation for the Riemannian metric $h_{u}$, where $u:=u(p)$.
Lemma 5.3.
Let $\gamma$ be a geodesic tangent to $\mathfrak{F}$, such that $\gamma(0)=p\in M$. Let $(u,v,x)$ be a system of coordinates around $p$ as above. If $x(t)$ is defined on $[0,\alpha]$, then $\gamma(t)$ is defined on $[0,\alpha]$.
Proof.
From the equation on $v(t)$, one can see that $v(t)$ is defined as soon as $x(t)$ is defined. However, given some coordinate chart $I\times J\times B_{n}(0,1)$, the $v(t)$-coordinate may well leave the interval $I$ before $x(t)$ leaves the ball $B_{n}(0,1)$. Suppose this is the case. One can slide over the kernel of $\mathfrak{F}$ and consider a new coordinate chart containing the geodesic leaving the first one. Remember that the flow of any vector field tangent to the kernel is isometric with respect to the Riemannian metric induced in the $\mathfrak{F}$-leaf containing $\gamma$. It follows that in this new chart, the equation on $x(t)$ is the same as in the first coordinate chart. One can do this until the $x$-coordinate leaves the ball $B_{n}(0,1)$. So $\gamma$ is indeed defined as soon as $x(t)$ is defined. ∎
Proof of Theorem 5.1.
[…]
5.2. Applications
As a consequence of Theorem 5.1, we can state a leafwise completeness result in the following situations
Corollary 5.4 (Leafwise completeness in presence of a parallel null line field).
Let $(M,l)$ be a compact Lorentzian manifold with a parallel null line field $l$. Then the leaves of the foliation tangent to the distribution $l^{\perp}$ are complete if and only if the geodesics tangent to $l$ are complete.
Corollary 5.5 (Leafwise completeness in presence of a null Killing field).
Let $(M,V)$ be a compact Lorentzian manifold with a null Killing field $V$, such that the distribution orthogonal to $V$ is integrable and defines a foliation $\mathfrak{F}$. Then the leaves of $\mathfrak{F}$ are complete.
6. On completeness of compact Lorentzian manifolds with a null Killing field
Let $(M,V,g)$ be a compact Lorentzian manifold with a null Killing field $V$. Recall that for any geodesic $\gamma$, we have $g(\dot{\gamma},V)=C$ constant along $\gamma$, and $C$ is called the Clairaut’s constant. In this section, we deal with the global completeness problem, i.e. completeness of the whole manifold $M$. We have the following result from the previous section
Theorem 6.1.
Let $(M,V)$ be a compact Lorentzian manifold with a null Killing field $V$, such that the distribution orthogonal to $V$ is integrable and defines a foliation $\mathfrak{F}$. Then the leaves of $\mathfrak{F}$ are complete.
We construct in this section an incomplete compact Lorentzian manifold with a null Killing field, which is locally homogeneous. We begin with some preliminaries on left invariant metrics on Lie groups.
6.1. Left invariant metrics on Lie groups
Let $G$ be a connected Lie group endowed with a left invariant pseudo-Riemannian metric, $i.e$ its Lie algebra ${\mathfrak{g}}$ is endowed with a non-degenerate inner product $q$. Given any curve $\gamma$ in $G$, we can consider the associated curve $y(t):=dL_{\gamma(t)}^{-1}(\dot{\gamma}(t))\in{\mathfrak{g}}$. By Euler-Arnold theorem, the geodesics of $G$ are in one-to-one correspondance with the integral curves of the (Euler-Arnold) quadratic-homogeneous vector field on ${\mathfrak{g}}$:
$y\in{\mathfrak{g}}\mapsto\dot{y}=\mathrm{ad}_{y}^{*}y\in{\mathfrak{g}}$ (1)
Null Killing fields
Let $\mathsf{v}\in{\mathfrak{g}}$ such that $\mathrm{ad}_{\mathsf{v}}$ is skew symmetric with respect to $q$. Then the left invariant metric on $G$ is also invariant by the right multiplication by $\exp(t\mathsf{v})$. So the left invariant vector field $V$ generated by $\mathsf{v}$ is Killing, and it is everywhere isotropic. Moreover, it acts isometrically (on the right) on $\Gamma\backslash G$, where $\Gamma\subset G$ is a cocompact lattice, defining a null Killing field on the compact quotient.
Completeness of geodesics orthogonal to $V$
It follows from the observation below that when $\mathsf{v}$ is null and $\mathrm{ad}_{\mathsf{v}}$ is skew symmetric, the left invariant distribution generated by the linear hyperplane $\mathsf{v}^{\perp}$ is lightlike and totally geodesic, and the geodesics tangent to it are exactly those orthogonal to the null Killing field $V$. We prove in Proposition 6.3 that those geodesics are complete (without assuming any compactness assumption).
Observation 6.2.
[…]
Proof.
Indeed, one has $g(\mathrm{ad}^{*}_{y}y,v)=g(\mathrm{ad}_{y}\mathsf{v},y)=-g(\mathrm{ad}_{\mathsf{v}}y,y)=0$. ∎
Proposition 6.3.
[…]
Proof.
Let $(\mathsf{v},\mathsf{x}=(\mathsf{x_{1}},..,\mathsf{x_{n}}),\mathsf{u})$ be a basis of ${\mathfrak{g}}$ such that $(\mathsf{x},\mathsf{v})$ generate $\mathsf{v}^{\perp}$. Let $y(t)$ be an integral curve of $\dot{y}$ tangent to $\mathsf{v}^{\perp}$, and write $y(t)=(v(t),x(t),u(t))$ in the above basis. The hyperplane $\mathsf{v}^{\perp}$ is lightlike with a transverse Riemannian quadratic form. Since the geodesic vector field is tangent to the levels of the quadratic form $q$, then on $\mathsf{v}^{\perp}$, it is tangent to levels of the form $\mathbb{S}^{n}\times\mathbb{R}$, where $\mathbb{R}$ is the direction of $\mathsf{v}$ and $\dim\mathsf{v}^{\perp}=n+1$. In particular, the $x$-component is bounded, hence complete. On the other hand, one can check that $\mathrm{ad}^{*}_{v}v=0$; as a consequence, Equation (1) on $y(t)$ yields an equation on $\dot{v}$ which is linear in $v$, hence $v(t)$ is also complete. ∎
Dynamics of $\dot{y}$
We saw that the integral curves of $\dot{y}$ on the linear hyperplane $\mathsf{v}^{\perp}$ correspond to geodesics orthogonal to $V$, which are complete. On the other hand, the integral curves of $\dot{y}$ on the affine translates of $\mathsf{v}^{\perp}$ correspond to the geodesics with Clairaut’s constant $C\neq 0$. After affine reparametrization, we can suppose $C=\pm 1$, so that the dynamics of the vector field $\dot{y}$ reduces to the ones on the two hyperplanes $\mathsf{v}^{\perp}\pm\mathsf{u}$, where $\mathsf{u}\in{\mathfrak{g}}$ is such that $q(\mathsf{u},\mathsf{v})=1$. The incompleteness obtained in the next paragraph occurs on those hyperplanes.
6.2. Incomplete examples with a null Killing field
Example 6.4 (An incomplete left invariant Lorentzian metric with a null Killing field).
Consider the Lie algebra ${\mathfrak{g}}={\mathfrak{aff}}(\mathbb{R})\oplus\mathbb{R}$. Let $(T,X,V)$ be a basis such that $[T,X]=X$. Let $q$ be the Lorentzian quadratic form on ${\mathfrak{g}}$ given by
$q=\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},$
in the basis $(T,X,V)$. The Euler-Arnold vector field in this case is
$w=\begin{pmatrix}t\\ x\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto ad^{*}_{w}w=\begin{pmatrix}-xv\\ 0\\ tv\end{pmatrix}$
So, we have the system of equations
$\begin{cases}\dot{t}=-xv\\ \dot{x}=0\\ \dot{v}=tv\end{cases}$
For $x=1$, we get that $\ddot{t}=-\dot{v}=-tv=t\dot{t}=(\frac{1}{2}t^{2})^{\prime}$. Hence, $\dot{t}=\frac{1}{2}t^{2}$. So $w(s)=(\frac{-2}{s},\frac{2}{s^{2}},1)^{T}$ is the velocity curve of an incomplete geodesic on the Lie group $G=\sf{Aff}(\mathbb{R})\times\mathbb{R}$ endowed with the left invariant metric coming from $q$.
Example 6.5 (An incomplete compact Lorentzian manifold with a null Killing field).
Let ${\mathfrak{g}}={\mathfrak{sol}}\oplus\mathbb{R}$. Let $(T,X,Y,V)$ be a basis such that $[T,X]=X,[T,Y]=-Y$, and where $V$ generates the center. Consider the Lorentzian quadratic form $q$ given by
$q=\begin{pmatrix}1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\end{pmatrix}$
in the basis $(V,T,X,Y)$. In this case the Arnold-Euler vector field is
$w=\begin{pmatrix}t\\ x\\ y\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto ad^{*}_{w}w=\begin{pmatrix}-xv+y^{2}\\ 0\\ ty\\ tv\end{pmatrix}$
In other words we get the following system
$\begin{cases}\dot{t}=-xv+y^{2}\\ \dot{x}=0\\ \dot{y}=ty\\ \dot{v}=tv\end{cases}$
We see that the geodesic field is tangent to the hyperplane $y=0$, which is then totally geodesic. On the other hand, this hyperplane is the subalgebra ${\mathfrak{aff}}(\mathbb{R})\oplus\mathbb{R}$ with the (incomplete) Lorentzian scalar product given in Example 6.4 above. Hence incompleteness. We obtain a compact example by taking a quotient by a cocompact lattice.
Remark 6.6.
In the previous example, the incomplete geodesics are timelike and lightlike.
7. Compact Lorentzian $3$-manifolds with an equicontinuous null Killing field
Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$, with $V$ a lightlike Killing field. In this section, we study the geodesic completeness of $M$. As pointed out in the introduction, the situation is different, depending on the dynamics of $V$. The non-equicontinuous case being already known, we focus here on the equicontinuous case. In this case, the closure of the flow of $V$ in the isometry group is a torus. Denote it by $G$.
Reading plan
An incomplete example is constructed in Paragraph 7.3.2, in the case where the flow of $V$ is periodic. The reader only needs Paragraph 7.2 to read it. Completeness results are given in special situations when the action of $V$ is not periodic, in paragraphs 7.3.3 and 7.3.1.
7.1. On the isometry group
We begin by a study of the connected component of the isometry group. We will see in particular that when it has dimension $\geq 3$, the manifold is flat.
Lemma 7.1.
Let $M$ be a connected compact Lorentzian $3$-manifold. If ${\sf{Isom}}(M)$ contains a $3$-dimensional connected abelian subgroup $H$, then $M$ is flat. Moreover, the identity component of the isometry group is isomorphic to $\mathbb{T}^{3}$.
Proof.
Assume first that $H$ is compact, i.e. $H=\mathbb{T}^{3}$. Let $K_{1},K_{2},K_{3}$ be three transversal Killing fields generated by the action of $H$. Let $C_{i}$ be the set of zeros of $K_{i}$, and let $C=\bigcup_{i=1}^{3}C_{i}$. Each of the $C_{i}$ has empty interior (this comes from the fact that the action of the isometry group is faithful). Consequently, $C$ is a closed subset of $M$, necessarily with empty interior (it is even of zero measure). So let $x\in M\smallsetminus C$. Then the $H$-orbit of $x$ is an open and closed subset of $M$, hence equal to $M$. This implies that the $H$-action is free and transitive, and $M$ is flat. Otherwise, $H$ is non-compact. By the construction above, there is a point $x_{0}\in M$ with a locally trivial isotropy. Hence, its orbit is a flat homogeneous open subset (of finite volume), we call it $U_{0}$. If $U_{0}$ is dense, then $M$ is flat, since the sectional curvature is zero on a dense subset of $M$. Otherwise, there is a point $x_{1}\in M\backslash U_{0}$, with also a locally trivial isotropy, and its orbit is a flat homogeneous orbit (of finite volume), we call it $U_{1}$. Moreover, $U_{0}\cap U_{1}=\emptyset$ and their volume is equal. This process must end for otherwise, $M$ would have infinite volume. When it ends, it means that there is some $U_{k}$ which is dense in the complement of $U_{0}\cup U_{1}\cup\cdots U_{k-1}$, so $M$ is covered by flat open subsets, hence flat. Now, since $M$ is flat (and compact), it is complete, and its isometry group is given by $\Gamma\backslash N_{\mathsf{Poi}^{1,2}}(\Gamma)$, where $N_{\mathsf{Poi}^{1,2}}(\Gamma)$ is the normalizer of $\Gamma:=\pi_{1}(M)$ in the Poincaré group. Since ${\sf{Isom}}(M)$ contains $H$, the normalizer of $\Gamma$ contains a $3$-dimensional (connected) abelian group, which is necessarily the translation group $\mathbb{R}^{3}$ by Observation 7.2 below. This group centralizes $\Gamma$, so $\Gamma$ is also made of pure translations (up to finite index). ∎
Observation 7.2.
The maximal dimension of an abelian connected subgroup of the Poincaré group ${\mathsf{Poi}}^{2,1}:={\sf{Isom}}({\sf{Mink}}^{1,2})$ is $3$. And it is identified with the pure translations.
Proposition 7.3.
Let $(M,V)$ be a connected compact Lorentzian $3$-manifold with an equicontinuous null Killing field $V$, with closure $G$ in ${\sf{Isom}}(M)$.
  1. (1)
    If $\dim G\geq 3$, then the identity component of the isometry group is isomorphic to $\mathbb{T}^{3}$, in which the flow of $V$ is dense, and $M$ is a flat $3$-torus.
  2. (2)
    If $\dim G=2$, then either $G$ coincides with the connected component of the isometry group, or, $M$ is flat and ${\sf{Isom}}^{\circ}(M)=\mathbb{T}^{3}$.
Proof.
$(1)$ By assumption, the closure of the flow of $V$ in ${\sf{Isom}}(M)$ is $G=\mathbb{T}^{k}$, with $k\geq 3$. The conclusion follows from Lemma 7.1.
$(2)$ Let ${\widehat{G}}:={\sf{Isom}}^{\circ}(M)$. First, recall that we have a bound on the dimension of ${\widehat{G}}$, namely $\dim{\widehat{G}}\leq 3+3$. The proof is based on the classification result on the connected component of the isometry group of a compact Lorentzian manifold [1, 2, 26].
Step 1: (Trivial non-compact part). Assume that the non-compact part of $\widehat{G}$ is non-trivial. We know via the classification result that the non-compact part (which, here, is either ${\sf{SL}}_{2}(\mathbb{R})\times\mathbb{R}^{k}$, ${\mathsf{Heis}}_{3}\times\mathbb{R}^{k}$, ${\mathsf{Osc}}\times\mathbb{R}^{k}$ or $\mathbb{R}^{k}$) commutes with the compact part, in particular with $\mathbb{T}^{2}$. Hence, ${\widehat{G}}$ contains a non-compact $3$-dimensional abelian subgroup given by the product of $\mathbb{T}^{2}$ with a one parameter group of the non-compact factor. This is absurd by Lemma 7.1.
Step 2: (Abelian compact part). Assume that the compact part of ${\widehat{G}}$ is non abelian. Since it contains $\mathbb{T}^{2}$, it is of the form $\mathbb{T}^{2}\times K$, where $K$ is, up to finite cover, isomorphic to either ${\sf{SO}}(3)$ or ${\sf{SO}}(3)\times\mathbb{S}^{1}$ (remember that $\dim K\leq 4$). Both cases will generate an abelian isotropy of dimension $\geq 2$, that represents (faithfully) in ${\sf{SO}}(1,2)$, which is absurd.
It follows from Step 1 and Step 2 that ${\widehat{G}}$ is compact abelian, isomorphic to $\mathbb{T}^{k}$, with $k\geq 2$. If $k=3$, $M$ is flat and ${\widehat{G}}=\mathbb{T}^{3}$ by Lemma 7.1. ∎
We refer to Paragraph 7.3.3 for examples where the identity component of the isometry group is a $2$-torus.
7.2. Existence of a $V$-invariant frame field
Lemma 7.4.
Let $(M,g,V)$ be an orientable compact Lorentzian manifold of dimension $3$, with $V$ a null Killing field. Then there is a null frame field $(V,X,Y)$ on $M$ which is invariant by the flow of $V$, and such that the metric in the basis $(V,Y,X)$ is given by $\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\\ \end{pmatrix}.$ Conversely, given a $V$-invariant frame field $(V,X,Y)$ in which the metric has the above form, $V$ is a null Killing field for this metric.
Proof.
The flow of $V$ preserves a Riemannian metric $h_{0}$ on $M$. Let $P_{0}$ (resp. $P$) be the $2$-dimensional distribution $h_{0}$-orthogonal ($g$-orthogonal) to $V$. Then $P_{0}\cap P$ is a line bundle over $M$. Define a vector field $X$ on $M$ tangent to the line bundle, and satisfying $g(X,X)=1$. Let $Y$ be the isotropic vector field on $M$ which is $g$-orthogonal to $X$ and satisfies $g(V,Y)=1$. Since the flow of $V$ is isometric for both $g$ and $h_{0}$, these vector fields are invariant by the flow of $V$, i.e. $[V,X]=[V,Y]=0$. For the converse, one proves that $\nabla V$ is skew-symmetric for $g$, using that the scalar products $g(U,W)$ are $V$-invariant for $U,W\in\{V,X,Y\}$. ∎
In particular, the foliation tangent to $V$ is trasnversally parallelizable. The action of $V$ is assumed to be equicontinuous. So the closure of the leaves of $V$ are the fibers of a fibration defined by the (proper) action of $G$ (the closure of the flow of $V$ in the isometry group). When the action of $V$ is free (observe that the action also preserves the orientation of $M$), the base space of this fibration is a compact and orientable manifold. Its Euler characteristic is zero, since the $V$-action preserves non-vanishing vector fields transversal to $V$, who then project to well defined vector fields on the base space. The base is therefore diffeomorphic to a torus (of dimension $\leq 2$), which we denote by $\Sigma$.
Lemma 7.5.
Let $(V,X,Y)$ be a null frame field as above. Then there exist $V$-invariant functions $f,h,\mu\in C^{\infty}(M,\mathbb{R})$ such that
  • $\nabla_{V}V=0$, $\nabla_{X}V=-\frac{1}{2}fV$, $\nabla_{Y}V=\frac{1}{2}fX$,
  • $[X,Y]=fY+hX+\mu V$.
In particular, $V$ is parallel if and only if $f=0$. Moreover, $f$, $h$ and $\mu$ are constant along the orbits of $V$, so they project to $\Sigma$.
Proof.
The first point is a straightforward computation, using that $V$ is Killing and all the scalar products involving $V,X,$ and $Y$ are constant. The fact that $f$, $h$ and $\mu$ are $V$-invariant follows from the fact that the frame is $V$-invariant, hence $[X,Y]$ is also $V$-invariant. ∎
7.3. Geodesic completeness
We begin now the investigation of the global completeness.
Observation 7.6.
In dimension $3$, the geodesics orthogonal to $V$ are complete. Indeed, the distribution orthogonal to $V$ is integrable, hence complete by Corollary 5.5.
7.3.1. Case of $\dim G\geq 3$
It appears from Proposition 7.3 that this case corresponds to a null Killing field $V$ having a dense orbit in $M$ (in which case, all orbits of $V$ are dense). We saw in Proposition 7.3 that in this case, $M$ is a flat $3$-torus, which is then in particular complete, and $\dim G=3$.
7.3.2. Case of $\dim G=1$
In this paragraph, we assume that $\dim G=1$, i.e. the flow of $V$ is periodic. Then $M$ is a Seifert fibration. For simplicity, we will assume that the $\mathbb{S}^{1}$-action of $V$ is free. Then $V$ is tangent to the fibres of a fibre bundle $\pi:M\to\Sigma$, where $\Sigma$ is diffeomorphic to a $2$-torus.
Getting started
The following lemma is a converse for Lemma 7.5 above.
Lemma 7.7.
Let $\pi:M\to\Sigma$ be an $\mathbb{S}^{1}$-bundle over a $2$-torus $\Sigma$, given by an $\mathbb{S}^{1}$ free action. Denote by $V$ the vector field on $M$ generated by the $\mathbb{S}^{1}$-action (it is tangent to the fibers). Let $(X^{*},Y^{*})$ be two vector fields on $\Sigma$ such that $[X^{*},Y^{*}]=f^{*}Y^{*}+h^{*}X^{*}$, with $f^{*},h^{*}\in C^{\infty}(\Sigma,\mathbb{R})$. There exist vector fields $(X,Y)$ on $M$ together with a Lorentzian metric $g$, such that
  • $(V,X,Y)$ is a $V$-invariant null frame field on $M$, and $V$ is a null Killing field for $g$,
  • $[X,Y]=fY+hX+\mu V$, with $f,h,\mu\in C^{\infty}(M,\mathbb{R})$ are $V$-invariant functions such that $f,h$ project to $f^{*},h^{*}$ respectively.
This metric is not unique.
Proof.
Choose a Riemannian metric $h_{0}$ on $M$ preserved by the flow of $V$. It defines a horizontal distribution on $M$ by taking the plane distribution orthogonal to $V$. Then $X$ and $Y$ are obtained by taking the horizontal lifts of $X^{*}$ and $Y^{*}$ respectively. By definition, $X$ and $Y$ project to $\Sigma$, so the coefficients $f,h,\mu$ are $V$-invariant. Define $g$ so that $(V,X,Y)$ is a null frame as in Lemma 7.4. The fact that $V$ is a Killing field follows from Lemma 7.4. ∎
Remark 7.8.
The distribution ${\sf{span}}(X,Y)$ is integrable if and only if $\mu=0$. In the previous lemma, one may consider a trivial fiber bundle $\Sigma\times\mathbb{S}^{1}$. In this case, since we have a global section, one can lift $X^{*}$ and $Y^{*}$ so that the distribution generated by $X$ and $Y$ is tangent to the sections, hence integrable.
Geodesic equation
Let $\gamma$ be a geodesic of $M$ and write
$\dot{\gamma}(t)=a(t)V+b(t)X+c(t)Y$
in the global framing. Let $\beta=g(\dot{\gamma},\dot{\gamma})$ and $\alpha=g(\dot{\gamma},V)$. Both $\beta$ and $\alpha$ are constant, and we have $c(t)=\alpha$ for any $t$.
Observation 7.9.
When $\alpha=0$, $\gamma$ is contained in a leaf of the $\mathfrak{F}$-foliation, and is then complete.
In the sequel, we consider geodesics transverse to the $\mathfrak{F}$-foliation, i.e. for which $\alpha\neq 0$.
Lemma 7.10 (Computation of Christoffel symbols).
The only non-null Christoffel symbols are
  • $\Gamma_{VY}^{X}=\Gamma_{YV}^{X}=-\Gamma_{VX}^{Y}=-\Gamma_{XV}^{Y}=\Gamma_{XY}^{V}=\frac{1}{2}f$
  • $\Gamma_{XY}^{X}=-\Gamma_{XX}^{Y}=h$
  • $\Gamma_{YY}^{X}=-\Gamma_{YX}^{Y}=\mu$
Proof.
Straightforward computation, using Lemma 7.5. ∎
So the geodesic equation when $\alpha\neq 0$ reads
$\dot{b}$ $=\frac{1}{2}f(\gamma(t))b^{2}-\alpha h(\gamma(t))b-\Big{(}\frac{\beta}{2}f(\gamma(t))+\alpha^{2}\mu(\gamma(t))\Big{)}$
$\dot{a}$ $=-\frac{f(\gamma(t))}{2\alpha}b^{3}+h(\gamma(t))b^{2}+\Big{(}\mu\alpha+\frac{\beta}{2\alpha}f(\gamma(t))\Big{)}b$
$\dot{c}$ $=0$
Reduction: Lorentzian metric associated to a triple $(\Sigma,X^{*},Y^{*})$
By Lemma 7.7 and Remark 7.8, one can associate to a triple $(\Sigma,X^{*},Y^{*})$ such that $[X^{*},Y^{*}]=f^{*}Y^{*}+h^{*}X^{*}$, a Lorentzian metric on the trivial bundle $\Sigma\times\mathbb{S}^{1}$ for which the lifts $X,Y$ of $X^{*}$ and $Y^{*}$ to $M$ define an integrable distribution. And one has $[X,Y]=fY+hX$, where $f$ and $h$ are fiber-constant functions obtained from $f^{*}$ and $h^{*}$ respectively.
Observation 7.11.
The orbits of $Y$ are totally geodesic. Indeed, we have $g(\nabla_{Y}Y,Y)=0$ and $g(\nabla_{Y}Y,V)=-g(Y,\nabla_{Y}V)=0$. And under the above reduction, i.e. integrability of the distribution ${\sf{span}}(X,Y)$, we have more: the flow of $Y$ is geodesic. This follows from $g(\nabla_{Y}Y,X)=\Gamma_{YY}^{X}=\mu=0$, which yields $\nabla_{Y}Y=0$. In this case, the orbits of $Y$ are complete null geodesics.
One asks:
Question 7.12.
What restrictions on $f^{*}$ and $h^{*}$ in order to obtain an incomplete example?
With the above reduction, the equation on $b(t)$ for an isotropic geodesic reads
$\dot{b}=\frac{1}{2}f(\gamma(t))b^{2}-\alpha h(\gamma(t))b.$ (2)
Observation 7.13.
If $b(t_{0})=0$ for some $t_{0}$, then $b$ is identically zero. In this case, $\gamma$ is a complete null geodesic everywhere tangent to ${\sf{span}}(Y,V)$.
Construction of an incomplete example
In what follows, the vector fields $X^{*}$ and $Y^{*}$ are lifted to $\widetilde{\Sigma}=\mathbb{R}^{2}$, keeping the same notations. So consider on $\mathbb{R}^{2}$ the two following linearly independent vector fields, written in the $(x,y)$-coordinates of the plane:
$X^{*}$ $=\cos(x)\partial_{x}-\sin(x)\partial_{y}$
$Y^{*}$ $=\sin(x)\partial_{x}+\cos(x)\partial_{y}$
By definition, they project to vector fields on a torus $\Sigma=\mathbb{T}^{2}$. We have
$[X^{*},Y^{*}]=\partial_{x}=\cos(x)X^{*}+\sin(x)Y^{*}.$
Consider the Lorentzian metric $(M,g,V)$ (with $V$ a null Killing field) on the trivial bundle $M=\Sigma\times\mathbb{S}^{1}$, associated to the triple $(\Sigma,X^{*},Y^{*})$. The two functions $f,h\in C^{\infty}(\Sigma,\mathbb{R})$ are given by $f(x,y)=\sin(x)$ and $h(x,y)=\cos(x)$.
The foliations determined by $X^{*}$ and $Y^{*}$ on $\widetilde{\Sigma}$ are illustrated in Figure 1.
[…]
Figure 1. The foliation tangent to $X^{*}$ (resp. to $Y^{*}$) is represented in black (resp. red)
They are made of Reeb components. This choice will allow Lemma 7.14 below, which is the key tool to incompleteness phenomena obtained in Corollary 7.15.
Let $\gamma$ be a null geodesic of $M$ defined on $I=[0,T[$, $T\in\overline{\mathbb{R}}$, such that $\alpha=1$, and $b(0)>0$. Denote by $\Gamma(t)=b(t)X^{*}+Y^{*}$ the projection of $\gamma$ to $\Sigma$ (lifted to $\mathbb{R}^{2}$). We introduce here some vocabulary we will use in the proof of Lemma 7.14; it is introduced to facilitate proof writing.
Vocabulary:
  • We call a band of $X^{*}$ (resp. of $Y^{*}$) any single Reeb component of $X^{*}$ (resp. of $Y^{*}$), with its boundary leaves.
  • We call half a band the intersection of a band of $X^{*}$ and a band of $Y^{*}$.
  • We say that $\Gamma$ crosses a band when there exist two distinct $t_{0},t_{1}\in I$ such that $\Gamma(t_{0})$ and $\Gamma(t_{1})$ are on distinct boundary leaves of the band. The same can be said in the case of a half band.
  • We say that $\Gamma$ makes a turnaround when there exist two distinct $t_{0},t_{1}\in I$ such that $\Gamma(t_{0})$ and $\Gamma(t_{1})$ belong to the boundary leaf of some band.
The boundary leaves of the bands coincide with the $x$-axis and its $\frac{k\pi}{2}$-translates, for $k\in\mathbb{Z}$.
Lemma 7.14.
Let $\gamma$ be a null geodesic of $M$ defined on $I=[0,T[$, $T\in\mathbb{R}\cup\{+\infty\}$, such that $\alpha=1$, and $b(0)>0$. Then
  1. (1)
    $\Gamma(t)$ does not cross any band.
  2. (2)
    $\Gamma(t)$ does not make turnarounds.
As a consequence, for $t$ close enough to $T$, $\Gamma(t)$ remains in half a band.
Proof.
(1) $\Gamma(t)$ does not cross any band. Indeed, if $\gamma$ crosses a band of $X$, then there is a point where $\gamma$ is tangent to $X$. But this means that $\alpha=0$. On the other hand, if $\gamma$ crosses a band of $Y$, then there is a point where $\gamma$ is tangent to $Y$. This means that $b(t_{0})=0$ for some $t_{0}\in I$. By Observation 7.13, this yields $b$ identically zero. In both cases, this contradicts our assumptions on $\gamma$.
(2) Indeed, a turnaround, when $\alpha\neq 0$ implies that $b$ vanishes. ∎
Corollary 7.15.
Let $l$ be the (boundary) leaf of $X^{*}$ such that $x(l)=\frac{\pi}{2}+2k\pi,k\in\mathbb{Z}$. Let $\gamma$ be a null geodesic of $M$ such that $p:=\Gamma(0)\in l$, $\alpha=1$, and $b(0)>0$. Then $\gamma$ is incomplete.
Proof.
By periodicity, we can fix $l$ to be the (boundary) leaf of $X$ such that $x(l)=\frac{\pi}{2}$. Let $l_{1}$ be the (boundary) leaf of $Y^{*}$ such that $x(l_{1})=\pi$. In $\mathbb{R}^{2}$, we can write $\Gamma(t)=(x(t),y(t))$, so that $\Gamma^{\prime}(t)=x^{\prime}(t)\partial_{x}+y^{\prime}(t)\partial_{y}$. On the other hand, $\Gamma^{\prime}(t)=b(t)X^{*}+Y^{*}$ (here we supposed $\alpha=1$). So, a straightforward computation yields
$x^{\prime}(t)$ $=b(t)\cos x+\sin x$ (3)
Step 1: We first prove that $\Gamma$ remains in the half band delimited by $l$ and $l_{1}$, so that $x(t)$ is bounded on $\Gamma$. The geodesic can leave the half band either by crossing the band, or by making a turnaround. Looking at the frame field $(X^{*},Y^{*})$ on the boundary leaves (see Figure 1), it appears that the first behaviour cannot occur. More precisely, if $\Gamma$ crosses the half band, then there is a point $t_{0}$ such that $x(t_{0})=\pi$ and $x^{\prime}(t_{0})\geq 0$. But Equation (3) yields $b(t_{0})\leq 0$, which is impossible. On the other hand, it follows from Lemma 7.14 that $\Gamma$ does not make turnarounds. The claim follows.
Observe now that in the interior of the half band, we have $f(x,y)=\sin x>0$ and $h(x,y)=\cos x<0$. So for $t$ close enough to 00 we get
$\dot{b}(t)$ $=\frac{1}{2}f(\Gamma(t))b^{2}(t)-h(\Gamma(t))b(t)$
$\geq\frac{1}{2}f(\Gamma(t))b^{2}(t),$
which is then actually true for any $t\in[0,T[$. In particular, $b(t)$ is an increasing function of $t$, and $\lim_{t\to T}b(t)=b_{\infty}\in\mathbb{R}_{>0}\cup\{+\infty\}$.
Step 2 ($f>0$ and bounded away from 00 on $\Gamma$): The way the frame field $(X,Y)$ varies in the half band forces the existence of some $\delta>0$ such that $\Gamma$ remains in the region $\frac{\pi}{2}<x<\pi-\delta$ for $t$ close enough to the limit $T$. Assume this is not the case. Then one can construct a sequence of points $t_{n}\to T$ such that $x(t_{n})$ goes to $\pi$. For such a sequence, we have $\lim_{n}x^{\prime}(t_{n})=-b_{\infty}\in\mathbb{R}_{<0}\cup\{-\infty\}$. If $x^{\prime}(t)$ was negative for $t$ close enough to the limit, meaning that $x(t)$ is decreasing, one cannot have $x(t_{n})$ converging to $\pi$ (since $x(t)<\pi$ for any $t\in[0,T[$). It follows that there is a sequence $t^{\prime}_{n}\to T$ such that $x^{\prime}(t^{\prime}_{n})=0$. But this means that $b(t)$ goes to 00, and this is a contradiction (remember that $b(t)$ is increasing and $b(0)>0$). The claim follows. There is no loss of generality in assuming that the claim holds for any $t\in[0,T[$. Then $f(\Gamma(t))>\epsilon$ for any $t\in[0,T[$, and some $\epsilon>0$.
Conclusion: Any solution of the differential equation $\dot{\tilde{b}}(t)=\frac{\epsilon}{2}\tilde{b}^{2}(t)$, with initial condition $\tilde{b}(0)>0$, is incomplete, defined on a bounded interval $[0,T_{0}[$, where $T_{0}<\infty$. For such a solution $\tilde{b}(t)$, we have $\lim_{t\to T_{0}}\tilde{b}(t)=\infty$, hence $\lim_{t\to T_{0}}\dot{\tilde{b}}(t)=\infty$. Since now $\dot{b}(t)\geq\dot{\tilde{b}}(t)$, incompleteness of $\Gamma$ with such an initial value $b(0)$ follows. ∎
Comment 7.16.
The general completeness question in dimension $3$ in the case of a periodic action reduces to Question 7.12. It would be interesting to investigate this question. It seems to us that the completeness depends on the dynamics of the vector fields $X^{*}$ and $Y^{*}$ defined on the torus.
7.3.3. Case of $\dim G=2$
This case corresponds to a non-periodic null Killing field $V$ with no dense orbit. The fact that there is a $\mathbb{T}^{2}$-action means that there is another Killing vector field $K$. Since $K$ is preserved by $V$, the norm of $K$ is constant on a $\mathbb{T}^{2}$-leaf, but it may change from a leaf to another; in particular, the type of $K$ may change. So the leaves of the $\mathbb{T}^{2}$-action may be either all Lorentzian (when $K$ is timelike for example), all degenerate, or both cases happening on the same manifold. There are special situations where one obtains completeness rather easily, but the general case is largely open.
Fact 7.17.
Let $(M,V)$ be a compact Lorentzian $3$-manifold with an equicontinuous null Killing field $V$, such that the closure of the flow of $V$ in ${\sf{Isom}}(M)$ is isomorphic to $\mathbb{T}^{2}$. Assume that the $\mathbb{T}^{2}$-action is free. Then
  • (1)
    If all the $\mathbb{T}^{2}$-leaves are degenerate, then the manifold is complete.
  • (2)
    If all the $\mathbb{T}^{2}$-leaves are Lorentzian, then the manifold is complete.
Proof.
(1) Let $K$ be a Killing field transversal to $V$, hence necessarily spacelike. Let $Y$ be the (unique) null vector field such that $g(Y,K)=0$ and $g(Y,V)=1$. Then $(V,K,Y)$ is a $V$-invariant frame field. Let $\gamma$ be a geodesic, and write $\dot{\gamma}(t)=a(t)V+b(t)K+c(t)Y$. We have that $c(t)=g(\dot{\gamma}(t),V)$ and $g(\dot{\gamma}(t),K)=b(t)g(K,K)$ are constant, and that $a(t)$ satisfies a differential equation $\dot{a}(t)=F(\gamma(t))$, where $F(p):=\alpha\frac{\nabla_{Y_{p}}g(K,K)}{g(K(p),K(p))^{2}}$, $\alpha\in\mathbb{R}$, is a bounded function on $M$ (see the note after Example 7.20, where we give the explicit computations leading to the equation on $a(t)$). Since $K$ is spacelike and $M$ is compact, the component $b(t)$ is a bounded function. Moreover, $a(t)$ is bounded on finite time intervals. Thus, the components of $\dot{\gamma}$ are bounded on finite time intervals, and the completeness follows.
(2) To prove completeness, we will show that there is a timelike Killing field generated by the action of $\mathbb{T}^{2}$. This action is free and proper, so the quotient map $\pi:M\to M/\mathbb{T}^{2}\simeq\mathbb{S}^{1}$ is a $\mathbb{T}^{2}$-principal bundle. Denote by $\mathcal{F}$ the $2$-dimensional foliation tangent to the fibers, and by $x$ the coordinate on the base space $\mathbb{S}^{1}$. Since the $\mathbb{T}^{2}$-action is isometric, it preserves the unitary vector field $T$ tangent to the orthogonal distribution $T\mathcal{F}^{\perp}$. It follows that the $\mathbb{T}^{2}$-action commutes with the flow of $T$. Denote by $C_{0}$ (resp. $C_{x}$) the nullcone field of the Lorentzian metric on $F_{0}$ (resp. on $F_{x}$). Let $x\in\mathbb{S}^{1}$, the (non oriented) angle $\alpha_{x}$ of the timecone delimited by $C_{x}$ is well-defined; indeed, since $\mathbb{T}^{2}$ is abelian, the Lorentzian metric on the fiber preserved by the $\mathbb{T}^{2}$-action is flat, so the nullcone field $C_{x}$ is parallel along the fiber. Moreover, since the parameter $x$ varies in a compact set, and the nullcone field varies continuously, $\alpha_{x}$ has a minimum $>0$. Now, to guarantee the existence of a timelike Killing field, we proceed as follows. For two leaves $F_{0}$ and $F_{x}$, $x\in\mathbb{S}^{1}$, there is a diffeomorphism $\phi^{t_{x}}_{T}:F_{0}\to F_{x}$, where $\phi^{t}_{T}$ is the flow of $T$. The pullback of the nullcone field on $F_{x}$ by the flow of $T$ is another nullcone field on $F_{0}$, whose (pullbacked) angle is $\alpha_{x}$. Since $M$ has a null Killing field $V$ commuting with $T$, the nullcone field $C_{0}$ and the pullbacks of the $C_{x}$, for all $x\in\mathbb{S}^{1}$, have a common (invariant) null vector field $V_{|F_{0}}$. This means that we can choose a timelike vector ${\mathfrak{z}}$ tangent to $F_{0}$ at some point $p\in F_{0}$, with the property that all its push-forwards by the $T$-flow are also timelike. Since the (isometric) action of $\mathbb{T}^{2}$ is transitive on the fibers, there exists a Killing field $K$ such that $K(p)={\mathfrak{z}}$. Since $K$ is $T$-invariant and all $d\phi_{T}^{t}({\mathfrak{z}})$ are timelike, $K$ is timelike along the orbit of $T$ through $x$ (which intersects all the fibers). Finally, the $\mathbb{T}^{2}$-invariance of $K$ proves that it is timelike everywhere. ∎
Remark 7.18.
In the proof of Item $(2)$ above, the flow of $T$ does not necessarily preserve the null cone fields of the leaves. In fact, those are preserved by the flow of $T$ if and only if the flow induces a conformal diffeomorphism between the leaves. The foliation $\mathcal{F}^{\perp}$ is totally geodesic (normal-(((which is equivalent to $\mathcal{F}$ being transversally Riemannian). And $\mathcal{F}$ is umbilic (which is equivalent to $\mathcal{F}^{\perp}$ being transversally conformal, see [24, Fact 2.1])normal-))), and since $\dim\mathcal{F}^{\perp}=1$, this is equivalent to the flow of $T$ preserving the conformal structure of the leaves of $\mathcal{F}$. When it is the case, by [24, Proposition 2.4], this leads to the warped product structure in Example 7.19 below.
Example 7.19 (Warped product).
Consider $\mathbb{R}^{3}$ with the Lorentzian metric $g_{f}=f(z)(dx^{2}-dy^{2})+dz^{2}$, where $f(z)>0$ and $l$-periodic. When $f$ is non-constant, $(\mathbb{R}^{3},g_{f})$ is non-flat. Moreover, the isometry group is, up to some discrete factor, $\mathbb{R}^{2}\times l\mathbb{Z}$. Hence, the quotient $M_{f}:=\mathbb{R}^{3}/\Gamma$ by a uniform lattice $\Gamma\subset\mathbb{R}^{3}$ is a compact Lorentzian manifold with connected component of the isometry group isomorphic to $\mathbb{T}^{2}$. The isometry group contains the flow of a null Killing field, and one can choose $\Gamma$ in such a way that this flow is dense in $\mathbb{T}^{2}$.
Example 7.20 (Non-warped product).
Consider $\mathbb{R}^{3}$ with the Lorentzian metric $g=(dudv+\cos(t)du^{2})+dt^{2}$. Let $V:=\partial_{v},U:=\partial_{u},T:=\partial_{t}$. Both $U$ and $V$ are Killing fields (with $V$ null) generating an $\mathbb{R}^{2}$-isometric action with Lorentzian flat leaves. The null cone field is spanned by $U$ and $U-\frac{1}{2}cos(t)V$. Since $[T,U-\frac{1}{2}\cos(t)V]=\frac{1}{2}\sin(t)V$ (normal-(((here, $[T,V]=0,[T,U]=0$)normal-))), the line field $\mathbb{R}(U-cos(t)V)$ is not preserved by $T$. So the flow of $T$ does not preserve the null cone field, and the decomposition $(\mathbb{R}^{2}\times\mathbb{R},(dudv+\cos(t)du^{2})+dt^{2})$ is therefore not a warped product. Note that this example is in fact a pp-wave, written in Brinkmann coordinates.
Note: Explicit computations leading to the geodesic equation in Fact 7.17
Let $(V,K,Y)$ be a $V$-invariant frame field such that $V$ and $K$ are Killing, and the scalar products $g(K,V),g(K,Y),g(V,V),g(Y,Y)$ and $g(Y,V)$ are constant. We claim that all Christoffel symbols $\Gamma_{**}^{Y}$ are null but $\Gamma_{KK}^{Y}=\frac{1}{2}\nabla_{Y}g(K,K)$. The only non-straightforward computation is $\Gamma_{KV}^{Y}$. Using that $K$ and $V$ are both Killing, that the scalar products $g(K,Y),g(Y,V)$ are constant, and that $[V,K]=0$, we compute the scalar products:
$\bullet\;\;g([K,Y],K)=g(\nabla_{K}Y,K)-g(\nabla_{Y}K,K)$ $=-g(\nabla_{K}K,Y)-g(\nabla_{Y}K,K)$
$=g(\nabla_{Y}K,K)-g(\nabla_{Y}K,K)=0$
$\bullet\;\;g([K,Y],Y)=g(\nabla_{K}Y,Y)-g(\nabla_{Y}K,Y)=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;$
$\bullet\;\;g([K,Y],V)=g(\nabla_{K}Y,V)-g(\nabla_{Y}K,V)$ $=-g(\nabla_{K}V,Y)+g(\nabla_{V}K,Y)$
$=g(\nabla_{V}K,Y)+g(\nabla_{V}K,Y)=0.$
This proves that $[K,Y]=0$. Using this, compute $g(\nabla_{Y}V,K)=-g(\nabla_{K}V,Y)=g(\nabla_{K}Y,V)=g(\nabla_{Y}K,V)=-g(\nabla_{Y}V,K)$, which proves that $g(\nabla_{Y}V,K)=0$. From this one infers that $\Gamma_{KV}^{Y}=\Gamma_{VK}^{Y}=0$. On the other hand, it is clear that $\Gamma_{KK}^{Y}=\frac{1}{2}\nabla_{Y}g(K,K)$. Finally, the vanishing of the other Christoffel symbols follow straightforwardly from the constancy of the scalar products in the lemma.